A certain spectral line in a specturm of a star has a frequencey 0.5x10^5 cyclic /Sec
If the star is approaching earth of a speed of 200 km by what te fraction will the frequency change?
1
Expert's answer
2019-11-18T11:24:29-0500

"f=\frac{c\pm{v}}{c}f_0=(1\pm{\frac{v}{c}})f_0"

"|f-f_0|=\frac{v}{c}f_0=\frac{200}{300000}*0.5*10^5=133.33\;cyclic/sec"

"\frac{|f-f_0|}{f_0}=\frac{133.33}{0.5*10^5}=0.0027"